# Bosonization in $R$-paraparticle Luttinger models

### v2

@@ 3 R-parafermion bosonization
$=\sum_{r,k}v_{F}(rk-k_{F})\Bigl{(}\hat{e}_{r,k,k}-m\theta(rk-k_{F})\Bigr{)}$ (14b)
where the constant $\theta(rk-k_{F})$ imposes the normal ordering denoted by $\ldots:$, and $\pm k_{F}$ are the parafermi points of the 1D system. The indices $k$ and $a$ denote momentum and flavor, while $r\in\{+,-\}$ (a mode index) differentiates between the two branches of the dispersion relation $\varepsilon_{r}(k)=v_{F}(rk-k_{F})$, with $v_{F}$ as the fermi velocity. In the Luttinger model, the total density operator can be decomposed in terms of the left-moving and right-moving $R$-parafermions, $\hat{\rho}(k)=\hat{\rho}_{+}(k)+\hat{\rho}_{-}(k)$, where
$\hat{\rho}_{r}(q)$ $=\sum_{k,a}\Bigl{(}\hat{\psi}^{+}_{r,k+q,a}\hat{\psi}_{r,k,a}-\delta_{q,0}\theta(rk-k_{F})\Bigr{)}$ (15a)
$=\sum_{k}\hat{e}_{r,k+q,q}-m\delta_{q,0}\sum_{k}\theta(rk-k_{F})$ (15b)
@@
$\hat{\sigma}_{r}(q)$ $=\sum_{k,a}\alpha_{a}\Bigl{(}\hat{\psi}^{+}_{r,k+q,a}\hat{\psi}_{r,k,a}-\delta_{q,0}\theta(rk-k_{F})\Bigl{)}.$ (15d)
For a direct comparison with the spin-$\frac{1}{2}$ Luttinger model, we restrict our results to the case of $m=2$ and define $\alpha_{1}=-1$ and $\alpha_{2}=+1$ when dealing with flavor waves. Note also that Eq. (15d) cannot be expressed in terms of contracted bilinear operators defined in Eq. (4) due to the prefactor $\alpha_{a}$.
3.1 Density wave bosonization and flavor-charge separation
Using Eq. (5), the commutators of the density operator can be evaluated as
$\hat{\rho}_{r}(-q),\hat{\rho}_{r^{\prime}}(q^{\prime})]=\delta_{rr^{\prime}}\delta_{qq^{\prime}}\sum_{k}(\hat{e}_{r,k-q,k-q}-\hat{e}_{r,k,k}),$ (16)
to which by invoking the properties of the parafermi points [28] and Eq. (11b), simplifies to
@@ 3.1 Density wave bosonization and flavor-charge separation
($L$ being the system length) dependent only on the maximum number of particles that can occupy a single mode. Comparison with the bose commutation relation $\hat{b}_{k},\hat{b}^{+}_{k}]=\delta_{kk^{\prime}}$ motivates the mapping
@@
$\hat{b}_{-q}$ $=\sqrt{\frac{2\pi}{n^{\prime}qL}}\hat{\rho}_{-}(q),\qquad\hat{b}_{-q}^{+}=\sqrt{\frac{2\pi}{n^{\prime}qL}}\hat{\rho}_{-}(-q),$
which then imply that the density waves of all $R$-parafermionic Luttinger models have bosonic properties.
Using the same commutation relations of the contracted bilinear operator, it can also be shown that
$\hat{\rho}_{r}(-q),\hat{\sigma}_{r^{\prime}}(q^{\prime})]=0.$ (19)
This means that for all $m=2$ cases, density waves and flavor waves are decoupled and propagate independently, thus implying that flavor-charge separation occurs for all $R$-parafermions.
3.2 Flavor wave bosonization
While Eq. (19) might suggest that both $\hat{\rho}_{r}(q)$ and $\hat{\sigma}_{r}(q)$ are in general bosonic operators of different species, explicit calculation of the commutation relations of the flavor wave operators show that it is not always the case. Because we cannot express $\hat{\sigma}_{r}(q)$ in terms of $\hat{e}_{ij}$, we are required to evaluate
@@ 3.2 Flavor wave bosonization
$=\sum_{k,k^{\prime},a,a^{\prime}}\alpha_{a}\alpha_{a^{\prime}}\Bigl{(}\hat{\psi}^{+}_{r,k-q,a}\hat{\psi}_{r,k,a}\hat{\psi}^{+}_{r^{\prime},k^{\prime}+q^{\prime},a^{\prime}}\hat{\psi}_{r^{\prime},k^{\prime},a^{\prime}}$
$\quad-\hat{\psi}^{+}_{r^{\prime},k^{\prime}+q^{\prime},a^{\prime}}\hat{\psi}_{r^{\prime},k^{\prime},a^{\prime}}\hat{\psi}^{+}_{r,k-q,a}\hat{\psi}_{r,k,a}\Bigr{)}$
$=\sum_{k,k^{\prime},a,a^{\prime}}\alpha_{a}\alpha_{a^{\prime}}\biggl{\{}\delta_{r,r^{\prime}}\delta_{k-q,k^{\prime}}\delta_{a,a^{\prime}}\hat{\psi}^{+}_{r,k-q,a}\hat{\psi}_{r^{\prime},k^{\prime},a^{\prime}}$
$\qquad+\biggl{(}\sum_{b,c,d,e,f,g}R^{ab}_{a^{\prime}c}R^{de}_{ab}R^{a^{\prime}c}_{gf}\hat{\psi}^{+}_{r^{\prime},k^{\prime}+q^{\prime},d}\hat{\psi}^{+}_{r,k-q,e}\hat{\psi}_{r^{\prime},k^{\prime},f}\hat{\psi}_{r,k,g}\biggr{)}$
$\qquad-\hat{\psi}^{+}_{r^{\prime},k^{\prime}+q^{\prime},a^{\prime}}\hat{\psi}_{r^{\prime},k^{\prime},a^{\prime}}\hat{\psi}^{+}_{r,k-q,a}\hat{\psi}_{r,k,a}\biggr{\}}$ (20)
explicitly for specific $R$-tensors. In this case, bosonic flavor waves occur only when the RHS of Eq. (3.2) reduces to the RHS of Eq. (16). This requirement is equivalent to
@@
as the second term in Eq. (21) cancels out the last term in Eq. (3.2) and the remaining Kronecker delta terms reduce to $\hat{e}_{ij}$. When the above requirement is satisfied, Eq. (3.2) becomes
@@
allows for the establishment of $\hat{c}_{k},\hat{c}^{+}_{k}]=\delta_{kk^{\prime}}$. $R$-tensors in the form of Eq. (13a), (13b), and (13d) with $\beta^{2}=1$ all follow these results. However, $R$-parafermions described by Eqs. (13c) and (13d) with $\beta^{2}\neq 1$ do not satisfy Eq. (21) and therefore their $R$-parafermionic Luttinger models do not have bosonic flavor wave profiles.
3.3 Equivalence spectrum
We now check wether the system described using $R$-parafermion operators are the same as that described by bosonic ones. For $m=2$ $R$-parafermions, it can be shown that the free Luttinger model Hamiltonian obeys the commutation relation
@@ 3.3 Equivalence spectrum
$\hat{H}_{0}=v_{F}\Bigl{[}\sum_{k}|k|\hat{n}_{k}+\frac{\pi}{2L}\sum_{r,a}(\hat{N}_{r,a})^{2}\Bigr{]}$ (25)
where $\hat{n}_{k}=\hat{b}_{k}^{+}\hat{b}_{k}$ and $\hat{N}_{r,a}=\sum_{k}(\hat{\psi}_{r,k,a}^{+}\hat{\psi}_{r,k,a}-\theta(rk-k_{F}))$. Note that these results are applicable only to $R$-parafermions with bosonic flavor waves. The second summation term takes care of the normal ordering imposed in the $R$-parafermion operator basis. The flavor wave contribution to the Hamiltonian is the same as Eq. (25), but with $\hat{b}_{k}^{+}\hat{b}_{k}$ replaced by $\hat{c}_{k}^{+}\hat{c}_{k}$ [29], hence we consider the total Hamiltonian
@@
where $H_{0}$ is denoted by Eqs. (14) and (25) in the $R$-parafermion and boson operator bases, respectively.
One way to verify the correspondence between these Hamiltonian forms is to show that their degeneracies match in both bases. The case where $R$-parafermions have no internal degrees of freedom (i.e., spinless fermions) has already been done in [26] by evaluating the partition functions of each basis separately. Following this method, the $R$-parafermion partition function is calculated as
$z_{\text{PF}}(y)=\prod_{n=1}^{\infty}\biggl{[}\Bigl{(}\sum_{j=0}^{p}\frac{1}{y^{2(2n-1)(j-1)}}\Bigr{)}\Bigl{(}\sum_{j=0}^{p}y^{2(2n-1)j}\Bigr{)}\biggr{]}^{4},$ (27)
where $y\equiv e^{-\beta v_{F}\pi/L}$. The (2n1)(2n-1) exponent is obtained by assuming that the system is under periodic boundary conditions such that $k=2n\pi/L$ (where $n$ are energy level indices) and $k_{F}=n_{F}\pi/L$. The variable $p$ denotes the order of the $R$-parafermion as discussed in Sec. 2.2. In particular, for $p=1$, we have
$z_{\text{PF}}(y)=\prod_{n=1}^{\infty}\biggl{[}1+y^{2(2n-1)}\biggr{]}^{8},$ (28a)
while for $p=2$,
$z_{\text{PF}}(y)=\prod_{n=1}^{\infty}\biggl{[}\Bigl{(}y^{2(2n-1)}+1+y^{-2(2n-1)}\Bigr{)}\Bigl{(}1+y^{2(2n-1)}+y^{4(2n-1)}\Bigr{)}\biggr{]}^{4}.$ (28b)
Meanwhile, the boson partition function is
$Z_{B}(y)=\biggl{[}\sum_{n=-\infty}^{+\infty}(y)^{n^{2}}\biggr{]}^{4}\prod_{n=1}^{\infty}\biggl{(}\frac{1}{1-y^{4n}}\biggr{)}^{4}=\prod_{n=1}^{\infty}\biggl{[}\frac{(1+y^{(2n-1)})^{2}(1-y^{2n})}{(1-y^{4n})}\biggr{]}^{4},$ (29)
the second equality being a result of the elliptic theta function equivalence
@@
Comparison between Eqs. (28a) and (29) shows that in the low-temperature limit (i.e., $\beta\to\infty$, and subsequently $y\to 0$), the boson and $p=1$ $R$-parafermion partition functions coincide, hence the spectrum of the free Luttinger model is equivalent in the $R$-paraparticle and boson operator basis only if the $R$-parafermion obeys Pauli’s exclusion principle. This result generalizes the idea that the Luttinger model is acceptable only at low-energies [46]. On the other hand, if the $R$-parafermion possesses a higher-order exclusion principle (such as in the case of $p=2$ in Eq. (28b)), $z_{PF}$ quickly diverges. This means that bosonization is not applicable for $R$-parafermions with $p>1$, and possibly the Luttinger model is not appropriate for modelling such systems.
4 Potential experimental realization
The Luttinger model assumes a linearized single-particle dispersion which makes it valid only for particles with momentum close to $\pm k_{F}$. In the fermionic Luttinger model literature, all ideal one-dimensional systems exhibit spin-charge separation, but for realistic materials these conditions are not always met and systems that satisfy the above requirements are said to reside in the Luttinger liquid phase. Several works have experimentally verified the existence of the Luttinger liquid phase by observing some of its key signatures [47, 48, 49, 50, 51, 31]. We highlight a potential experimental design for observing $R$-paraparticle signatures by borrowing ideas from [31], which we outline as follows.
In the above paper, Luttinger liquid signatures were detected as two parabolic dispersions due to separate fermi seas with different masses, which are associated with spin and charge excitations. Since $R$-parafermionic Luttinger models also exhibit flavor-charge separation, we should expect that a one-dimensional system containing $R$-parafermions would also display such behavior. Current predictions show that $R$-paraparticles can occur as quasiparticle excitations [23]. In this case, the fermi momentum in one dimension, expressed as [55]
$k_{F}=\frac{N_{0}\pi}{n^{\prime}L},$ (31)
is in general not equal for different $R$-parafermions. Here, $N_{0}$ denotes the number of particles occupying the ground state, and assuming that the system is made up of fundamental fermions which in turn have emergent $R$-parafermion quasiparticle properties, the number of fermions in the ground state is generally not equal to the number of $R$-parafermions in the quasiparticle ground state. By then constructing a system which does not exhibit spin-density excitations (e.g., there is no spin-charge separation or the system is fully spin-polarized) but contains emergent $R$-paraparticles that are in the Luttinger liquid phase, then signatures of flavor-charge separation may also be detected as separate dispersion relations as in [31].
The above proposal is roughly sketched, but it does give us a roadmap of research extensions to undertake in order to achieve an experimental observation of $R$-paraparticle signatures through Luttinger liquids. Firstly, a method which can turn a system of ordinary fermions or bosons to quasiparticle $R$-parafermion excitation is required. In [23], it was shown that emergent one-dimensional quasiparticle excitations with parastatistical GCRs can be developed from a generic spin model where the spin operators have $R$-paraparticle behaviors. For example, emergent $R$-paraparticles described by Eq. (13a) were shown to be realizable using three-level Rydberg atom or molecule systems [56, 57], while those described by Eq. (13b) with $\alpha=1$ can be modelled using systems described in [58]. In general, several papers have proposed a method of mapping between particles of different statistics in one-dimensional systems [59, 60], although an extension to the $R$-paraparticles is yet to be explored.
Nonlinear extensions of the fermionic Luttinger model [61, 46, 62] have also shown that spin-charge separation can persist outside the previously predicted low-energy regime [31]. While this may allow for easier detection of flavor wave oscillations at higher temperatures, it may also induce spin-charge separation at the same time. We are then required to find parameter regimes where spin-wave excitations cease to exist while flavor waves persist.
Further complications to this proposal would be due to the limitations of our results to $R$-paraparticles with $2$ internal degrees of freedom. The construction of the flavor wave operator was inspired using an analog of spin waves in spin-$\frac{1}{2}$ systems. While quantum spins are not directly observable, electron spin-wave propagation can be observed experimentally [42]. As such, Eq. (15d) should be sufficient to describe a wave propagation due to some internal degree of freedom of an $R$-paraparticle with $m=2$. This allows us to potentially observe $R$-paraparticle signatures even without full knowledge of the internal symmetries at play within the system. However, if the emergent $R$-paraparticle contains $m>2$ internal degrees of freedom, then the flavor wave operator must be redefined to account for multiple types of flavor wave oscillations, while at the same time, preserving the total degrees of freedom of the system.
5 Summary and Discussion

### v3

@@ III R-parafermion bosonization
$=\sum_{r,k}v_{F}(rk-k_{F})\Bigl(\hat{e}_{r,k,k}-n^{\prime}\theta(k_{F}-rk)\Bigr)$ (15b)
where the constant $\theta(rk-k_{F})$ imposes the normal ordering denoted by $\ldots:$, and $\pm k_{F}$ are the parafermi points of the 1D system. The indices $k$ and $a$ denote momentum and flavor, while $r\in\{+,-\}$ (a mode index) differentiates between the two branches of the dispersion relation $\varepsilon_{r}(k)=v_{F}(rk-k_{F})$, with $v_{F}$ as the fermi velocity. In the LM, the total density operator can be decomposed in terms of the left-moving and right-moving $R$-parafermions, $\hat{\rho}(k)=\hat{\rho}_{+}(k)+\hat{\rho}_{-}(k)$, where
$\hat{\rho}_{r}(q)$ $=\sum_{k,a}\Bigl(\hat{\psi}^{+}_{r,k+q,a}\hat{\psi}_{r,k,a}-\delta_{q,0}\theta(k_{F}-rk)\Bigr)$ (16a)
$=\sum_{k}\hat{e}_{r,k+q,q}-n^{\prime}\delta_{q,0}\sum_{k}\theta(k_{F}-rk)$ (16b)
@@
$\hat{\sigma}_{r}(q)$ $=\sum_{k,a}\alpha_{a}\Bigl(\hat{\psi}^{+}_{r,k+q,a}\hat{\psi}_{r,k,a}-\delta_{q,0}\theta(k_{F}-rk)\Bigl).$ (16d)
For a direct comparison with the spin-$\frac{1}{2}$ LM, we restrict our results to the case of $m=2$ and define $\alpha_{1}=-1$ and $\alpha_{2}=+1$ when dealing with flavor waves. Note also that Eq. (16d) cannot be expressed in terms of contracted bilinear operators defined in Eq. (4) due to the prefactor $\alpha_{a}$.
III.1 Density wave bosonization and flavor-charge separation
Writing the density operator in terms of $\hat{e}_{r,i,j}$ allows us to evaluate its commutation relation using Eq. (5), resulting to
$\hat{\rho}_{r}(-q),\hat{\rho}_{r^{\prime}}(q^{\prime})]=\delta_{rr^{\prime}}\delta_{qq^{\prime}}\sum_{k}(\hat{e}_{r,k-q,k-q}-\hat{e}_{r,k,k}).$ (17)
We then note that for $R$-parafermions, we can exploit their fermi-surface structures which leads to the closed form expression [39]
@@ III.1 Density wave bosonization and flavor-charge separation
($L$ being the system length) dependent only on the maximum number of particles that can occupy a single mode. Comparison with the bose commutation relation $\hat{b}_{k},\hat{b}^{+}_{k}]=\delta_{kk^{\prime}}$ then motivates the mapping
@@
$\hat{b}_{-q}$ $=\sqrt{\frac{2\pi}{n^{\prime}qL}}\hat{\rho}_{-}(q),\qquad\hat{b}_{-q}^{+}=\sqrt{\frac{2\pi}{n^{\prime}qL}}\hat{\rho}_{-}(-q).$
Note that because Eq. (5) applies to all $R$-tensors, the above equation implies that all $R$-parafermionic LMs possess density waves with bosonic properties. This includes $R$-parafermion types with more than two internal degrees of freedom, and is a property irrespective of order.
Using the same commutation relations of the contracted bilinear operators, it can also be shown that
$\hat{\rho}_{r}(-q),\hat{\sigma}_{r^{\prime}}(q^{\prime})]=0,$ (20)
which is applicable for all $R$-parafermions with $m=2$. However, to interpret this commutation relation as flavor-charge separation, we must also show that the flavor-wave operator satisfies bose statistics.
III.2 Flavor wave bosonization
While Eq. (20) shows that $\hat{\rho}_{r}(q)$ and $\hat{\sigma}_{r}(q)$ always commute for $m=2$ $R$-parafermions, explicit calculation of the commutation relations of the flavor wave operators shows that it is not always bosonic in nature, hence, Eq. (20) does not fully capture flavor-charge separation in that sense. Because we cannot express $\hat{\sigma}_{r}(q)$ in terms of $\hat{e}_{ij}$, we are required to evaluate the commutator explicitly as
@@ III.2 Flavor wave bosonization
$=\sum_{k,k^{\prime},a,a^{\prime}}\alpha_{a}\alpha_{a^{\prime}}\biggl\{\delta_{r,r^{\prime}}\delta_{k-q,k^{\prime}}\delta_{a,a^{\prime}}\hat{\psi}^{+}_{r,k-q,a}\hat{\psi}_{r^{\prime},k^{\prime},a^{\prime}}-\hat{\psi}^{+}_{r^{\prime},k^{\prime}+q^{\prime},a^{\prime}}\hat{\psi}_{r^{\prime},k^{\prime},a^{\prime}}\hat{\psi}^{+}_{r,k-q,a}\hat{\psi}_{r,k,a}$
$\qquad+\biggl(\sum_{b,c,d,e,f,g}R^{ab}_{a^{\prime}c}R^{de}_{ab}R^{a^{\prime}c}_{gf}\hat{\psi}^{+}_{r^{\prime},k^{\prime}+q^{\prime},d}\hat{\psi}^{+}_{r,k-q,e}\hat{\psi}_{r^{\prime},k^{\prime},f}\hat{\psi}_{r,k,g}\biggr)\biggr\}$ (21)
showing us that the system bosonizes only if the RHS of Eq. (III.2) simplifies to an expression similar to Eq (17). This can only be achieved whenever the last term of the above expression reduces to
@@
or more simply, when the $R$-tensors satisfy
$\sum_{c,d}R^{ac}_{bd}R^{ef}_{ac}R^{bd}_{hg}=\delta_{e,b}\delta_{h,a}R^{ef}_{hg}.$ (23)
This allows for the reduction of the commutator to a closed form
@@
allows for the establishment of $\hat{c}_{k},\hat{c}^{+}_{k}]=\delta_{kk^{\prime}}$.
Returning to our earlier $M$-matrix examples, $M_{1}$, $M_{2}(\gamma\in\mathbb{C}_{\neq 0})$, and $M_{4}(\delta^{2}=1)$ satisfy Eq. (23) and hence must have bosonic flavor waves. However, $M_{3}$ and $M_{4}(\delta^{2}\neq 1)$ do not follow this requirement and therefore do not possess bosonic flavor wave profiles. We note that even though $M_{2}(\gamma=1)$ can be mapped to $M_{4}(\delta\in\mathbb{C}_{\neq 0})$ under internal basis transformations, the latter case can only satisfy Eq. (23) when $\delta=\pm 1$. This shows that Eq. (23) is a sufficient but not a necessary condition for flavor-wave bosonization, as it can still be satisfied by certain $R$-parafermions after some suitable internal basis transformation. The existence of $R$-parafermions that do not satisfy Eq. (23) for any of its internal bases remain an open possibility.
III.3 Equivalence spectrum
We now check wether the system described using $R$-parafermion operators are the same as that described by bosonic ones. For $m=2$ $R$-parafermions, it can be shown that the free LM Hamiltonian obeys the commutation relation
@@ III.3 Equivalence spectrum
$\hat{H}_{0}=v_{F}\Bigl[\sum_{k}|k|\hat{n}^{(b)}_{k}+\frac{\pi}{2L}\sum_{r,a}(\hat{N}_{r,a})^{2}\Bigr]$ (27)
where $\hat{n}^{(b)}_{k}=\hat{b}_{k}^{+}\hat{b}_{k}$ and $\hat{N}_{r,a}=\sum_{k}(\hat{\psi}_{r,k,a}^{+}\hat{\psi}_{r,k,a}-\theta(k_{F}-rk))$. The second summation term takes care of the normal ordering imposed in the $R$-parafermion operator basis. The flavor wave contribution to the Hamiltonian is the same as Eq. (27), but with $\hat{n}^{(b)}=\hat{b}_{k}^{+}\hat{b}_{k}$ replaced by $\hat{n}^{(c)}=\hat{c}_{k}^{+}\hat{c}_{k}$ [40]. Note that these results are applicable only to $R$-parafermions with bosonic flavor waves. The total Hamiltonian under consideration would then be
@@
where $H_{0}$ is given by Eqs. (15b) and (27) in the $R$-parafermion and boson operator bases, respectively.
Figure 2: Dispersion relation of $R$-parafermionic LM with (a) contact, (b) Yukawa, and (c) screened interactions, where in (b) and (c), we used $g=1$. We have set $v_{F}=1$ for all calculations. For all cases, we see distinct differences between the flavor-wave (denoted by the broken line) and density-wave dispersions (in solid lines). The corresponding density wave velocity is obtained as the slope of the solid curves at $q=0$.
One way to verify the correspondence between these Hamiltonian forms is to show that their degeneracies match in both bases. The case where $R$-parafermions have no internal degrees of freedom (i.e., spinless fermions) has already been done in [37] by evaluating the partition functions of each basis separately. Following this method, we make use of the orthonormal $R$-parafermion Fock basis where
$\hat{e}_{r,k,k}|n_{r,k}\rangle$ $=\hat{n}_{r,k}|n_{r,k}\rangle=n_{r,k}|n_{r,k}\rangle,$ (29a)
$\langle n_{r,k}|n^{\prime}_{r^{\prime},k^{\prime}}\rangle$ $=\delta_{nn^{\prime}}\delta_{rr^{\prime}}\delta_{kk^{\prime}}.$ (29b)
By imposing the latter requirement, we are then restricted to unitary $R$-tensors since non-unitary $R$ requires a more involved inner product definition. The $R$-parafermion partition function is calculated as
$Z_{\text{PF}}(y)=\prod_{n=1}^{\infty}\biggl(\sum_{j=0}^{n^{\prime}}y^{2(2n-1)j}\biggr)^{4},$ (30)
where $y\equiv e^{-\beta v_{F}\pi/L}$. Note that this expression is independent of the order $p$ of the $R$-parafermion. The (2n1)(2n-1) exponent is obtained by assuming that the system is under periodic boundary conditions such that $k=2n\pi/L$ (where $n$ are energy level indices) and $k_{F}=n_{F}\pi/L$. Meanwhile, the boson partition function is obtained from the orthonormal basis set $\{|N_{r,a},n_{k}^{(b)},n^{(c)}_{q}\rangle\}$, where
$\hat{N}_{r,a}|N_{r,a},n_{k}^{(b)},n^{(c)}_{q}\rangle$ $=N_{r,a}|N_{r,a},n_{k}^{(b)},n^{(c)}_{q}\rangle$, (31a)
$\hat{n}^{(b)}_{k}|N_{r,a},n_{k}^{(b)},n^{(c)}_{q}\rangle$ $=n^{(b)}_{k}|N_{r,a},n_{k}^{(b)},n^{(c)}_{q}\rangle$, (31b)
$\hat{n}^{(c)}_{q}|N_{r,a},n_{k}^{(b)},n^{(c)}_{q}\rangle$ $=n^{(c)}_{q}|N_{r,a},n_{k}^{(b)},n^{(c)}_{q}\rangle$. (31c)
These are constructed from the boson vacuum state via
$|N_{r,a},n_{k}^{(b)},n^{(c)}_{q}\rangle=(\hat{U}_{r,a})^{N_{r,a}}\frac{(\hat{b}^{+}_{k})^{n_{k}^{(b)}}}{\sqrt{n_{k}^{(b)}!}}\frac{(\hat{c}^{+}_{q})^{n_{q}^{(c)}}}{\sqrt{n_{q}^{(c)}!}}|0\rangle,$ (32)
where $\hat{U}_{r,a}$ is the ladder operator for the $R$-parafermion charge $N_{r,a}$. This comes from a direct generalization of the ladder operators introduced in [37], and results to the boson partition function
$Z_{B}(y)$ $=\biggl[\sum_{n=-\infty}^{+\infty}(y)^{n^{2}}\biggr]^{4}\prod_{n=1}^{\infty}\biggl(\frac{1}{1-y^{4n}}\biggr)^{2}$
$=\prod_{n=1}^{\infty}\biggl[\frac{(1+y^{(2n-1)})^{4}(1-y^{2n})^{2}}{(1-y^{4n})}\biggr]^{2},$ (33)
where the second equality results from the elliptic theta function equivalence
@@
Fig 1(b) shows a comparison of the partition functions of the LM for both bases. It is interesting to note that $R$-parafermions with different $n^{\prime}$ have almost equal values of $Z_{PF}$ even at high temperatures. In contrast to the spinless LM, the boson and $R$-parafermion partition functions do not exactly match. However, $Z_{B}$ and $Z_{PF}$ shows a significant overlap in the low temperature limit, $\beta\to\infty$. This signifies that the spectra described by the Hamiltonians in the boson and $R$-parafermion bases are equivalent, thus implying that bosonization becomes a valid process for 1D systems at low temperatures, generalizing the existing predictions for fermionic LMs.
III.4 Interacting systems
In preparation for the next section, we now consider a more realistic system with inter-particle interaction. For simplicity, we make use of the total Hamiltonian
$\hat{H}^{LM}$ $=\hat{H}_{0}+\hat{H}_{FW}+\hat{H}_{int},$ (35a)
$\hat{H}_{int}$ $=\frac{1}{2L}\sum_{k}V(k)\hat{\rho}(k)\hat{\rho}(-k).$ (35b)
This can be solved exactly by transforming the Hamiltonian into the boson basis and then applying appropriate diagonalization techniques. The interaction term can be simplified using the mapping given by Eq. (19), while the free-particle term transforms as Eq. (27). However, under consideration that we conserve the particle number, then we can drop the second term of Eq. (27) [38], and the remaining problem becomes equuivalent to a spin-$\frac{1}{2}$ LM, where the total Hamiltonian can be decomposed in terms of density and flavor contributions,
$\hat{H}=\sum_{k}E_{\rho}(k)\hat{b}_{k}^{+}\hat{b}_{k}+E_{\sigma}(k)\hat{c}_{k}^{+}\hat{c}_{k},$ (36)
with corresponding dispersions [39]
$E_{\rho}(k)$ $=v_{F}|k|\sqrt{1+\frac{2V(k)}{v_{F}\pi}},$ (37a)
and
$E_{\sigma}(k)$ $=v_{F}|k|.$ (37b)
In Figs. 2(a)-(c), we plot the dispersion relations of interactions of the forms
$V(q)$ $=g,$ (38a)
$V(q)$ $=\frac{g}{q^{2}+\lambda^{-2}},$ (38b)
and
$V(q)$ $=ge^{-\Lambda|q|},$ (38c)
respectively, where the former potential corresponds to a repulsive contact interaction term, while the latter two are medium-ranged interactions, all for $g>0$. This shows that for a simple interacting $R$-parafermionic 1D system described by the LM, one can observe flavor-charge separation as two distinct dispersions with different velocity profiles. Finally, the partition function for the total interacting system is
$Z=\prod_{k=-\infty}^{\infty}\frac{1}{(1-e^{-\beta E_{\rho}(k)})(1-e^{-\beta v_{F}|k|})},$ (39)
and thermodynamic variables such as heat capacity and entropy can be calculated after specifying the system’s interaction potential.
IV Potential experimental realization
The LM assumes a linearized single-particle dispersion, which makes it valid only in the momentum regime close to $\pm k_{F}$. This idealization implies that not all real systems possess the necessary properties to exhibit LM properties such as spin-charge separation. Systems that realize these properties are therefore termed Luttinger liquids (LLs). Several works have experimentally verified the existence of LLs by observing some of its key signatures [54, 55, 56, 57, 58, 42]. Specifically in Ref. [42], they were able to verify spin-charge separation in a 1D system by looking for two distinct dispersions with different velocity profiles, which were associated with the decoupled spin and charge excitations. Since we showed that $R$-parafermionic LMs also exhibit flavor-charge separation, we should expect that a 1D system containing $R$-parafermions would also display such behavior.
Following Ref. [32], suppose that we were able to construct a system of emergent $R$-paraparticles from an underlying 1D fermion system with total particle number $N$. The mapping introduced in the latter ensures that the total particle number is conserved. By writing the fermi momentum as
$k_{F}=\frac{N\pi}{n^{\prime}L},$ (40)
we can deduce that the validity of the LL theory for the fermion system and the $R$-parafermion quasiparticles overlap if $n^{\prime}$ are equal for both cases. This may then be problematic, as it becomes difficult to distinguish when the separate dispersions are due to the underlying fermions or to the emergent quasiparticle system. This would be the case for $M_{2}(\gamma=1)$ $R$-paraparticle systems in the form of $m$ decoupled chains of $XY$ models which maps to free fermion chains [59].
In contrast, if the underlying system were bosonic, then we can disregard the issue of fermi momentum matching and possible dispersion overlaps. Consider for example a 1D spinor bose gas in the TG limit. In terms of the Bose-Hubbard model, we can write the Hamiltonian as
$\hat{H}=-t\sum_{i,a}(\hat{b}_{i,a}^{+}\hat{b}_{i+1,a}+\hat{b}_{i+1,a}^{+}\hat{b}_{i,a})-\sum_{i}\mu_{i}\hat{n}_{i},$ (41)
where $t$ is the hopping parameter and $\mu_{i}$ is the lattice chemical potential. The hard-core condition implies $n_{i}=\{0,1\}$, which imposes the restriction
$\hat{b}_{i,a}^{+}\hat{b}_{i,a^{\prime}}^{+}=0,\quad\hat{b}_{i,a}\hat{b}_{i,a^{\prime}}^{+}=\delta_{aa^{\prime}}\Bigl(1-\sum_{c}\hat{b}_{i,c}^{+}\hat{b}_{i,c}\Bigr)$ (42)
for local operators, while still having $\hat{b}_{i,a},\hat{b}_{j,a^{\prime}}^{+}]=[\hat{b}_{i,a},\hat{b}_{j,a^{\prime}}]=0$ for $i\neq j$ [60, 61]. This allows the bosons to be completely mapped to a system of emergent $R$-paraparticles described by $M_{1}$, with the Hamiltonian now defined as
$\hat{H}=-t\sum_{i,a}(\hat{\psi}_{i,a}^{+}\hat{\psi}_{i+1,a}+\hat{\psi}_{i+1,a}^{+}\hat{\psi}_{i,a})-\sum_{i}\mu_{i}\hat{n}_{i}.$ (43)
Reexpressing the Hamiltonian in momentum basis and expanding the kinetic energy term about $\pm k_{F}$ while taking $\mu_{i}=0$ allows us to obtain an $R$-paraparticle LM with $v_{F}=2t\sin k_{F}$. By including interaction terms in the underlying Hubbard model and mapping them appropriately, we can also introduce certain forms of $R$-paraparticle interactions in our system.
The important thing with spinor bose gases is that while they also exhibit spin-charge separation, the contribution of their spin components to the total energy is small due to a large effective mass in their dispersion. Specifically, in the TG limit, the effective mass becomes infinite and the spins become stationary [48]. We expect no dispersion profile due to boson spins for an $R$-paraparticle system from an underlying TG gas, and therefore flavor wave dispersions should be more easily identifiable, as sketched in Fig. 1(c). Owing these to properties, spinor TG gases should therefore be ideal candidates for the realization of flavor-charge separation in $R$-parafermion systems.
V Summary and Discussion